# Notes on Deterministic and Stochastic Approaches in Electromagnetic Information Theory

### v1

@@ 2 Some Recalls on the Non-Stochastic Approach to EIT
where $G(\mathbf{r},\mathbf{r}^{\prime})$ is the Green’s function in free space, equal to $\frac{e^{-j\beta R}}{4\pi R}$, $R=|\mathbf{r}-\mathbf{r}^{\prime}|$, $\mathbf{r}\in S_{obs}$, $\beta$ is the free-space wavenumber. ${E}^{d}$ and ${J}^{d}$ are square integrable functions and ${J}^{d}$ is bounded.
@@
These basis diagonalize the radiation operator, such that $y^{d}_{n}=\sigma_{h}x^{d}_{k}$. The functions ${u}_{n}(\mathbf{r}^{\prime})$ and ${v}_{n}(\mathbf{r}^{\prime})$ are respectively the left and right singular functions and $\sigma_{n}$ are the singular values of the operator [21]. Singular values are real non negative quantities. We suppose that they are sorted in descending way, i.e. $\sigma_{1}\geq\sigma_{2}\geq$.... The number of significant singular values fixes the Number of Degrees of Fredom ($N_{DoF}\,\,$) of the field [16]. The singular functions $v_{n}(\mathbf{r})$ and $u_{n}(\mathbf{r})$ are obtained as solution of eigenvalue problems [22]. In particular, $u_{n}(\mathbf{r})$ are the eigenfunctions of the eigenproblem whose kernel is:
@@
This approach, developed from studies on electromagnetic field representation [16, 17], makes the extensive literature on field interpolation directly applicable to EIT  [^1] For mathematical rigor, some care must be taken when moving from approximation theory to information theory. The former essentially involves a covering problem for the set of fields, while the latter requires solving a packing problem [5]. This results in a different $N_{DoF}\,\,$. However, this difference arises from the exclusion of only one singular value of the radiation operator and is therefore generally negligible..
@@ 3.1 Information carried by stochastic fields
Now, with reference to the same model shown in Fig. 1, let us consider a statistical ensemble of current distributions $\{j^{(r)}(\mathbf{r}^{\prime},t)\}$, where each realization $j^{(r)}$ represents a Cartesian component of the current density vector at a point specified by a position vector $\mathbf{r}^{\prime}\in D$ at time $t$. This ensemble is modeled as a stationary stochastic process, at least in the wide sense.
Analogously, we consider a statistical ensemble of radiated fields $\{e^{(r)}(\mathbf{r},t)\}$, stationary, at least in the wide sense, whose single realization is a Cartesian component of the electric field vector at a point specified by a position vector $\mathbf{r}\in S_{obs}$ at time $t$.
@@
The cross correlation function (called also mutual coherence) of the field current density in the source point $\mathbf{r}^{\prime}_{1}$ and $\mathbf{r}^{\prime}_{2}$ and the cross correlation function of the field in the observation points $\mathbf{r}_{1}$ and $\mathbf{r}_{2}$ and $\tau$ time delay interval, are [20](Sect. 4.2):
$\Gamma(\mathbf{r}_{1},\mathbf{r}_{2},\tau)$ $=\mathbb{E}\{{\hat{E}}(\mathbf{r}_{1},t){\hat{E}}^{*}(\mathbf{r}_{2},t+\tau)\}$ (5)
$\Gamma_{J}(\mathbf{r^{\prime}}_{1},\mathbf{r^{\prime}}_{2},\tau)$ $=\mathbb{E}\{{\hat{J}}(\mathbf{r^{\prime}}_{1},t){\hat{J}}^{*}(\mathbf{r^{\prime}}_{2},t+\tau)\}.$ (6)
where $\hat{E}$ is the complex analytic representation of $e^{r}(t,\mathbf{r})$ whose imaginary part is related to the original waveform by Hilbert transformation, $\mathbb{E}\{\hat{E}\}$ denotes the ensemble average of $\hat{E}$, and $\hat{E}^{*}$ is the complex conjugate of $\hat{E}$ [20](Sect. 3). A similar symbology is used for $\hat{J}$.
The discussion on the stochastic meaning of the field is analogous to that of the currents. For this reason, we’ll focus our discussion mainly on the field.
In $\hat{E}(\mathbf{r},t)$ each member is a possible realization of the electric field at every point in space and at every instant in time. This ensemble accounts for all the temporal fluctuations and spatial variations of the field.
The cross-spectral density function (CSD) is given by the generalized Fourier  [^2] For more details on generalized harmonic analysis see [20](Sect. 2.4). frequency transform of $\Gamma(\mathbf{r}_{1},\mathbf{r}_{2},t)$ (for currents $\Gamma_{J}(\mathbf{r}_{1},\mathbf{r}_{2},t)$) [20](Sect. 4.3):
$W(\mathbf{r}_{1},\mathbf{r}_{2},\omega)$ $=\frac{1}{2\pi}\int_{-\infty}^{\infty}\Gamma(\mathbf{r}_{1},\mathbf{r}_{2},t)e^{j\omega t}dt$ (7)
$W_{J}(\mathbf{r}^{\prime}_{1},\mathbf{r}^{\prime}_{2},\omega)$ $=\frac{1}{2\pi}\int_{-\infty}^{\infty}\Gamma_{J}(\mathbf{r}^{\prime}_{1},\mathbf{r}^{\prime}_{2},t)e^{j\omega t}dt$ (8)
In the following, anytime we will analyze the problem at a specific angular frequency $\omega=\bar{\omega}$, the dependence on $\omega$ will be understood and dropped, paralleling the harmonic signal representation. Accordingly, we will call $W(\mathbf{r}_{1},\mathbf{r}_{2})=W(\mathbf{r}_{1},\mathbf{r}_{2},\bar{\omega})$ and $W_{J}(\mathbf{r}_{1},\mathbf{r}_{2})=W_{J}(\mathbf{r}_{1},\mathbf{r}_{2},\bar{\omega})$
@@
where $\nabla^{2}_{1}$ and $\nabla^{2}_{2}$ operate respectively on $\mathbf{r}_{1}$ and on $\mathbf{r}_{2}$ keeping $\mathbf{r}_{2}$ and $\mathbf{r}_{1}$ fixed, and $\beta=\bar{\omega}/c$, $c$ being the velocity of the light in vacuum. Loosely speaking, cross-spectral density ’propagates’ in the space according to laws fixed by Maxwell’s equation.
The eigenfunctions $\psi_{n}(\mathbf{r})$ represent completely coherent modes. In general, $E(\mathbf{r})$ is only partially coherent. From its CSD we introduce the spectral degree of coherence [20](Sect. 4.3):
$\mu(\mathbf{r}_{1},\mathbf{r}_{2})=\frac{W(\mathbf{r}_{1},\mathbf{r}_{2})}{\sqrt{W(\mathbf{r}_{1},\mathbf{r}_{1})\,W(\mathbf{r}_{2},\mathbf{r}_{2})}}$ (16)
where $W(\mathbf{r}_{1},\mathbf{r}_{1})$ (and similarly $W(\mathbf{r}_{2},\mathbf{r}_{2})$) represents the spectral intensity of the field at point $\mathbf{r}_{1}$, denoted as $I(\mathbf{r}_{1})$ (and similarly $I(\mathbf{r}_{2})$), i.e., the average power per unit area per unit frequency at that specific location. Determining mutual information generally requires knowledge of the full joint probability distribution of the variables, which implies knowing the complete statistics of the underlying process. However, in the case of zero-mean stationary complex Gaussian circularly symmetric processes, [^3] The assumption of a Gaussian distribution is widely accepted in wireless communication systems, as it maximizes Shannon entropy and information content of the field. A circularly symmetric complex Gaussian process is characterized by a probability distribution that is invariant under phase rotations. This assumption implies that the pseudo-covariance matrix (which describes correlations between non-conjugate complex variables) is identically null, thereby simplifying the process’s characterization significantly. If the process has a non-zero mean, the CSD used for mutual information calculation must be evaluated on the fluctuating part of the stochastic process. That is, instead of $E(\mathbf{r}_{1})$ and $E(\mathbf{r}_{2})$, the CSD is evaluated on $E^{\prime}(\mathbf{r}_{1})=E(\mathbf{r}_{1})-\mathbb{E}\{E(\mathbf{r}_{1})\}$ and $E^{\prime}(\mathbf{r}_{2})=E(\mathbf{r}_{2})-\mathbb{E}\{E(\mathbf{r}_{2})\}$ [23]. the statistical properties are entirely described by covariance (or correlation) function, and the differential mutual information $I_{MI}(\mathbf{r}_{1};\mathbf{r}_{2})$ between the two points is given by [24]:
@@
In the following analysis, we will assume a spatially uncorrelated source  [^5] Rigorously speaking, a source cannot be perfectly spatially uncorrelated [25]. However, the correlation is subject to the low-pass filtering process of the free-space Green’s function that tends to filter the higher spatial frequencies [20] (Sect. 5.3). This implies that the approximation of a spatially incoherent source does not significantly affect the field results when the source’s intrinsic spatial correlation length is sufficiently short (let us say, shorter than $\lambda/2$), as propagation naturally increases the coherence distance. , which can be considered the ’richest source’ in terms of spatial information. For such a source, the cross-spectral density of the current is given by $W_{J}(\mathbf{r}^{\prime}_{1},\mathbf{r}^{\prime}_{2})=I(\mathbf{r}^{\prime}_{1})\delta(\mathbf{r}^{\prime}_{1}-\mathbf{r}^{\prime}_{2})$, where $I(\mathbf{r}^{\prime})=W_{J}(\mathbf{r}^{\prime},\mathbf{r}^{\prime})$ is the spectral intensity distribution of the source at point $\mathbf{r}^{\prime}$.
@@
$W(\mathbf{r}_{1},\mathbf{r}_{2})=\int_{D}I(\mathbf{r}^{\prime})G(\mathbf{r}_{1},\mathbf{r}^{\prime})G^{*}(\mathbf{r}_{2},\mathbf{r}^{\prime})d\mathbf{r}^{\prime}$ (19)
@@
In particular, the Van Cittert-Zernike theorem [20] (Sects. 4.4.5 and 5.5.1) states that under the Fraunhofer approximation, the CSD of a spatially incoherent planar source is equal to the Fourier transform of its intensity distribution. This result resembles the planar near-field to far-field transformation in antenna theory, and indeed shares many common characteristics. Consequently, the Van Cittert-Zernike theorem serves as a powerful tool to connect information and electromagnetism within the Fraunhofer zone.
The $N_{DoF}\,\,$of a stochastic process can be obtained by the Karhunen-Loève expansion. With reference to Eq. (9), we recall that the eigenvalues $\lambda_{n}$ are real and non-negative, and exhibit a step-like behavior: a finite number of them are significant (close to a maximum value), while the rest are negligible. The number of these significant modes is the $N_{DoF}\,\,$of the stochastic field. It can be easily shown that the $N_{DoF}\,\,$of the field radiated by an uncorrelated source coincides with the $N_{DoF}\,\,$of the non statistical model (1) provided that we consider a "weighted" radiation operator whose kernel is $G^{\prime}(\mathbf{r},\mathbf{r}^{\prime})=G(\mathbf{r},\mathbf{r}^{\prime})\sqrt{I(\mathbf{r}^{\prime})}$.
Let us consider the case an uncorrelated homogeneous  [^6] A homogeneous source is a source whose statistical properties are uniform across its surface, i.e. whose spectral density is constant on the source plane. source with $I(\mathbf{r}^{\prime})$=1. In this case the eigenproblem in Eq. (9) is the same eigenproblem that gives the left singular functions of the radiation operator, whose kernel is Eq. (4). Consequently, the eigenfunctions of the CSD operator coincide with the left singular functions of the radiation operator, and the eigenvalues of the CSD operator coincide with the square of the singular values of the radiation operator, giving a direct connection in terms of $N_{DoF}\,\,$between the stochastic and non-stochastic model.
This key equivalence explains the remarkable success of classical electromagnetic methods in the EIT analysis of spatial information.
@@ 3.2 $N_{DoF}\,\,$in stochastic field
In essence, for each spatial point $\mathbf{r}$, the Wigner function $\mathcal{W}(\mathbf{r},\mathbf{k})$ provides a "spectrum" of wavenumbers $\mathbf{k}$ that characterize the field’s coherence properties at that location. The marginal of $\mathcal{W}$ with respect to $\mathbf{k}$ yields the spatial distribution of the spectral intensity, while, the marginal with respect to position $\mathbf{r}$ yields the angular power spectrum. Finally integrating the WDF over the entire phase space (both $\mathbf{r}$ and $\mathbf{k}$) gives the total power $\Phi_{\text{total}}$ [28].
@@
where with Volume we mean the volume of Wigner function support in phase space, and $n$ represents the number of spatial dimensions being considered for the field’s transverse extent (i.e., the dimensionality of the spatial domain $\mathbf{r}$)  [^11] For example, $n=1$ in case of 1D spatial problem (e.g., a line source), $n=2$ if considering a 2D transverse plane (e.g., a planar source or observation screen). Since for each such spatial dimension, there is a corresponding conjugate wave vector dimension, the phase space itself is $2n$-dimensional. . For example, in case of a line source $\ell$ long the $N_{DoF}\,\,$in the $z>0$ half-plane ($n=1$) is given by the the volume of the phase space, $2\beta\times\ell$, divided by $2\pi$ 121212 Note the similarity of this approach to the Landau theorem in Information Theory [31] to the identification of $N_{DoF}\,\,$. As an example, in case of a band-limited complex signal having total (bilateral) angular bandwidth $\hat{\Omega}_{bil}$ acquired over an interval $T$, the $N_{DoF}\,\,$, derived from the Landau theorem, is given by the time-bandwidth product $\hat{\Omega}_{bil}\times T$ divided by $2\pi$..
@@
These results coincides with the results obtained in EIT using non-stochastcic approaches based on modal expansion [34], the Weyl’s law [35], the classic near field-far field transformation [33], [14], and with the results based on field sampling theory [17], apart from a $\pi/4$ factor  [18]  [^14] This discrepancy arises from the common sampling strategy in current electromagnetic field measurement systems, where sampling positions are constrained to lie along the coordinate curves of the scanning surface due to mechanical limitations [36]..
@@
This method becomes an exact method to evaluate the $N_{DoF}\,\,$in the high-frequency asymptotic regime ($\lambda\rightarrow 0$), i.e., the geometrical optics approximation. In this regime, the $N_{DoF}\,\,$is strictly related to the étendue ($G_{e}$) [37], a pure geometrical quantity of great importance in radiometry [20](Sect. 5.7)  [^15] Other interesting connections between radiometry and the stochastic approach used in this paper, that are not discussed here due to space limitations, are reported in [38]. since it is conserved under the geometrical optics approximation in the absence of diffraction, absorption, or scattering phenomena  [^16] This invariance corresponds to the conservation of entropy in the absence of absorption, emission, or scattering processes [39]..
@@
Figure 3: From a ’spatial information’ point of view, linear source sees the scene in 2 distorted.

### v3

@@ 2 Some Recalls on the Non-Stochastic Approach to EIT
where the free–space Green’s function is $G(\mathbf{r},\mathbf{r}^{\prime})=-\,j\bar{\omega}\mu_{0}\frac{e^{-j\beta R}}{4\pi R},\qquad R=|\mathbf{r}-\mathbf{r}^{\prime}|,\quad\mathbf{r}\in S_{\mathrm{obs}}$, $\beta=\bar{\omega}\sqrt{\epsilon_{0}\mu_{0}}$ is the free–space wavenumber, and $\epsilon_{0}$ and $\mu_{0}$ are the permittivity and permeability of free space, respectively. The functions $E^{d}$ and $J^{d}$ are square–integrable, and $J^{d}$ is bounded.
@@
These basis diagonalize the radiation operator, such that $y^{d}_{n}=\sigma_{h}x^{d}_{n}$. The functions ${u}_{n}(\mathbf{r}^{\prime})$ and ${v}_{n}(\mathbf{r}^{\prime})$ are respectively the left and right singular functions and $\sigma_{n}$ are the singular values of the operator [21]. Singular values are real non negative quantities. We suppose that they are sorted in descending way, i.e. $\sigma_{1}\geq\sigma_{2}\geq$.... The number of significant singular values fixes the Number of Degrees of Fredom ($N_{DoF}\,\,$) of the field [16]. The singular functions $v_{n}(\mathbf{r})$ and $u_{n}(\mathbf{r})$ are obtained as solution of eigenvalue problems [22]. In particular, $u_{n}(\mathbf{r})$ are the eigenfunctions of the eigenproblem whose kernel is:
@@
This approach, developed from studies on electromagnetic field representation [16, 17], makes the extensive literature on field interpolation directly applicable to EIT  [^1] For mathematical rigor, some care must be taken when moving from approximation theory to information theory. The former essentially involves a covering problem for the set of fields, while the latter requires solving a packing problem [5]. This results in a different $N_{DoF}\,\,$. However, this difference is therefore generally negligible for large $N_{DoF}\,\,$..
@@ 3.1 Information carried by stochastic fields
Now, with reference to the same model shown in Fig. 1, let us consider a statistical ensemble of current distributions $\{j^{(r)}(\mathbf{r}^{\prime},t)\}$, where each realization $j^{(r)}$ represents the current density at a point specified by a position vector $\mathbf{r}^{\prime}\in D$ at time $t$. This ensemble is modeled as a stationary stochastic process, at least in the wide sense.
Analogously, we consider a statistical ensemble of radiated fields $\{e^{(r)}(\mathbf{r},t)\}$, stationary, at least in the wide sense, whose single realization is the electric field at a point specified by a position vector $\mathbf{r}\in S_{obs}$ at time $t$.
@@
The cross-correlation function (also called mutual coherence) of the field at the observation points $\mathbf{r}_{1}$ and $\mathbf{r}_{2}$ with time delay $\tau$, and the cross-correlation function of the current density at the source points $\mathbf{r}^{\prime}_{1}$ and $\mathbf{r}^{\prime}_{2}$, are [20, Sec. 4.2]:
$\Gamma(\mathbf{r}_{1},\mathbf{r}_{2};\tau)$ $=\mathbb{E}\!\left\{\hat{E}(\mathbf{r}_{1},t)\,\hat{E}^{*}(\mathbf{r}_{2},t+\tau)\right\},$ (5)
$\Gamma_{J}(\mathbf{r}^{\prime}_{1},\mathbf{r}^{\prime}_{2};\tau)$ $=\mathbb{E}\!\left\{\hat{J}(\mathbf{r}^{\prime}_{1},t)\,\hat{J}^{*}(\mathbf{r}^{\prime}_{2},t+\tau)\right\}.$ (6)
where $\hat{E}$ is the complex analytic representation of $e^{r}(t,\mathbf{r})$ whose imaginary part is related to the original waveform by Hilbert transformation, $\mathbb{E}\{\hat{E}\}$ denotes the ensemble average of $\hat{E}$, and $\hat{E}^{*}$ is the complex conjugate of $\hat{E}$ [20](Sect. 3). A similar notation is used for $\hat{J}$.
The discussion on the stochastic meaning of the field is analogous to that of the currents. For this reason, we will focus mainly on the field.
In the ensemble $\{\hat{E}(\mathbf{r},t)\}$, each member is a possible realization of the electric field at every point in space and at every instant in time. This ensemble accounts for all temporal fluctuations and spatial variations of the field.
Under the wide-sense stationarity (WSS) assumption for both the field and the current, the temporal correlation functions depend only on the time lag $\tau$.
The cross-spectral density (CSD) is the Fourier transform in $\tau$ [20, Sec. 4.3]:
$W(\mathbf{r}_{1},\mathbf{r}_{2},\omega)$ $=\frac{1}{2\pi}\int_{-\infty}^{\infty}\Gamma(\mathbf{r}_{1},\mathbf{r}_{2};\tau)\,e^{+j\omega\tau}\,d\tau,$ (7)
$W_{J}(\mathbf{r}^{\prime}_{1},\mathbf{r}^{\prime}_{2},\omega)$ $=\frac{1}{2\pi}\int_{-\infty}^{\infty}\Gamma_{J}(\mathbf{r}^{\prime}_{1},\mathbf{r}^{\prime}_{2};\tau)\,e^{+j\omega\tau}\,d\tau.$ (8)
In the following, whenever we analyze the problem at a specific angular frequency $\omega=\bar{\omega}$, the dependence on $\omega$ will be understood and dropped, paralleling the harmonic signal representation. Accordingly, we will call $W(\mathbf{r}_{1},\mathbf{r}_{2})=W(\mathbf{r}_{1},\mathbf{r}_{2},\bar{\omega})$ and $W_{J}(\mathbf{r^{\prime}}_{1},\mathbf{r^{\prime}}_{2})=W_{J}(\mathbf{r^{\prime}}_{1},\mathbf{r^{\prime}}_{2},\bar{\omega})$
@@
where $\nabla^{2}_{1}$ and $\nabla^{2}_{2}$ operate respectively on $\mathbf{r}_{1}$ and on $\mathbf{r}_{2}$ keeping $\mathbf{r}_{2}$ and $\mathbf{r}_{1}$ fixed. Loosely speaking, cross-spectral density ’propagates’ in the space according to laws fixed by Maxwell’s equation.
The eigenfunctions $\psi_{n}(\mathbf{r})$ represent completely coherent modes. In general, $E(\mathbf{r})$ is only partially coherent. From its CSD we introduce the spectral degree of coherence [20, Sec. 4.3]:
$\mu(\mathbf{r}_{1},\mathbf{r}_{2})\;=\;\frac{W(\mathbf{r}_{1},\mathbf{r}_{2})}{\sqrt{\,W(\mathbf{r}_{1},\mathbf{r}_{1})\,W(\mathbf{r}_{2},\mathbf{r}_{2})\,}}$, (16)
where $W(\mathbf{r}_{1},\mathbf{r}_{1})$ (and similarly $W(\mathbf{r}_{2},\mathbf{r}_{2})$) is the spectral density of the electric field at the point $\mathbf{r}_{1}$, denoted $S(\mathbf{r}_{1})$ (and $S(\mathbf{r}_{2})$).
Determining mutual information generally requires knowledge of the full joint probability distribution of the variables, which implies knowing the complete statistics of the underlying process. However, in the case of zero-mean stationary complex Gaussian circularly symmetric processes, [^2] The assumption of a Gaussian distribution is widely accepted in wireless communication systems, as it maximizes Shannon entropy and information content of the field. A circularly symmetric complex Gaussian process is characterized by a probability distribution that is invariant under phase rotations. This assumption implies that the pseudo-covariance matrix (which describes correlations between non-conjugate complex variables) is identically null, thereby simplifying the process’s characterization significantly. If the process has a non-zero mean, the CSD used for mutual information calculation must be evaluated on the fluctuating part of the stochastic process. That is, instead of $E(\mathbf{r}_{1})$ and $E(\mathbf{r}_{2})$, the CSD is evaluated on $E^{\prime}(\mathbf{r}_{1})=E(\mathbf{r}_{1})-\mathbb{E}\{E(\mathbf{r}_{1})\}$ and $E^{\prime}(\mathbf{r}_{2})=E(\mathbf{r}_{2})-\mathbb{E}\{E(\mathbf{r}_{2})\}$ [23]. the statistical properties are entirely described by covariance (or correlation) function, and the differential mutual information $I_{MI}(\mathbf{r}_{1};\mathbf{r}_{2})$ between the two points is given by [24]:
@@
In the following analysis, we will assume a spatially uncorrelated source  [^4] Rigorously speaking, a source cannot be perfectly spatially uncorrelated [25]. However, the correlation is subject to the low-pass filtering process of the free-space Green’s function that tends to filter the higher spatial frequencies [20] (Sect. 5.3). This implies that the approximation of a spatially incoherent source does not significantly affect the field results when the source’s intrinsic spatial correlation length is sufficiently short (let us say, shorter than $\lambda/2$), as propagation naturally increases the coherence distance. , which can be considered the ’richest source’ in terms of spatial information. For such a source, the cross-spectral density of the current is given by $W_{J}(\mathbf{r}^{\prime}_{1},\mathbf{r}^{\prime}_{2})=S_{J}(\mathbf{r}^{\prime}_{1})\delta(\mathbf{r}^{\prime}_{1}-\mathbf{r}^{\prime}_{2})$, where $S_{J}(\mathbf{r}^{\prime}_{1})=W_{J}(\mathbf{r}^{\prime},\mathbf{r}^{\prime})$ is the source spectral density per unit angular frequency..
@@
$W(\mathbf{r}_{1},\mathbf{r}_{2})=\int_{D}S_{J}(\mathbf{r}^{\prime})G(\mathbf{r}_{1},\mathbf{r}^{\prime})G^{*}(\mathbf{r}_{2},\mathbf{r}^{\prime})d\mathbf{r}^{\prime}$ (19)
@@
In particular, the Van Cittert-Zernike theorem [20] (Sects. 4.4.5 and 5.5.1) states that under the Fraunhofer approximation, the CSD of a spatially incoherent planar source is proportional to the Fourier transform of its intensity distribution. This result resembles the planar near-field to far-field transformation in antenna theory, and indeed shares many common characteristics. Consequently, the Van Cittert-Zernike theorem serves as a powerful tool to connect information and electromagnetism within the Fraunhofer zone.
The $N_{DoF}\,\,$of a stochastic process can be obtained by the Karhunen-Loève expansion. With reference to Eq. (9), we recall that the eigenvalues $\lambda_{n}$ are real and non-negative, and exhibit a step-like behavior: a finite number of them are significant (close to a maximum value), while the rest are negligible. The number of these significant modes is the $N_{DoF}\,\,$of the stochastic field. It can be easily shown that the $N_{DoF}\,\,$of the field radiated by an uncorrelated source coincides with the $N_{DoF}\,\,$of the non statistical model (1) provided that we consider a "weighted" radiation operator whose kernel is $G^{\prime}(\mathbf{r},\mathbf{r}^{\prime})=G(\mathbf{r},\mathbf{r}^{\prime})\sqrt{S_{J}(\mathbf{r}^{\prime})}$.
Let us consider the case an uncorrelated homogeneous  [^5] A homogeneous source is a source whose statistical properties are uniform across its surface, i.e. whose spectral density is constant on the source. source with $S_{J}(\mathbf{r}^{\prime}_{1})$=1. In this case the eigenproblem in Eq. (9) is the same eigenproblem that gives the left singular functions of the radiation operator, whose kernel is Eq. (4). Consequently, the eigenfunctions of the CSD operator coincide with the left singular functions of the radiation operator, and the eigenvalues of the CSD operator coincide with the square of the singular values of the radiation operator, giving a direct connection in terms of $N_{DoF}\,\,$between the stochastic and non-stochastic model.
This key equivalence has a rather general validity (see Appendix A for a proof in the general case) and explains the remarkable success of classical electromagnetic methods in the EIT analysis of spatial information.
@@ 3.2 $N_{DoF}\,\,$in stochastic field
In essence, for each spatial point $\mathbf{r}$, the Wigner function $\mathcal{W}(\mathbf{r},\mathbf{k})$ provides a "spectrum" of wavenumbers $\mathbf{k}$ that characterize the field’s coherence properties at that location. The marginal of $\mathcal{W}$ with respect to $\mathbf{k}$ yields the spatial distribution of the spectral intensity, while, the marginal with respect to position $\mathbf{r}$ yields the angular spectrum. Finally, in a predominantly radiative field configuration, integrating the WDF over the propagating part of phase space (both $\mathbf{r}$ and $\mathbf{k}$) yields a quantity proportional to the radiative spectral power [^8] Under a local plane-wave (i.e., far-field) assumption, an electric-field–based Wigner description is sufficient to estimate power density. In that case the proportionality factor is $1/\!\big(2\,\eta_{0}\,(2\pi)^{n}\big)$, where $\eta_{0}$ is the free space wave impedance , $n$ is the dimensionality of the spatial domain $\mathbf{r}$, and $(2\pi)^{n}$ follows from the Fourier-transform normalization. This proportionality does not generally hold in more complex field configurations (e.g., in the reactive near field)..
@@
where with Volume we mean the volume of Wigner function support in phase space, and $n$ represents the number of spatial dimensions being considered for the field’s transverse extent (i.e., the dimensionality of the spatial domain $\mathbf{r}$)  [^11] For example, $n=1$ in case of 1D spatial problem (e.g., a line source), $n=2$ if considering a 2D transverse plane (e.g., a planar source or observation screen). Since for each such spatial dimension, there is a corresponding conjugate wave vector dimension, the phase space itself is $2n$-dimensional. . For example, in case of a line source $\ell$ long the $N_{DoF}\,\,$in the $z>0$ half-plane ($n=1$) is given by the the volume of the phase space, $2\beta\times\ell$, divided by $2\pi$  [^12] Note the similarity of this approach to the Landau theorem in Information Theory [30] to the identification of $N_{DoF}\,\,$. As an example, in case of a band-limited complex signal having total (bilateral) angular bandwidth $\hat{\Omega}_{bil}$ acquired over an interval $T$, the $N_{DoF}\,\,$, derived from the Landau theorem, is given by the time-bandwidth product $\hat{\Omega}_{bil}\times T$ divided by $2\pi$..
@@
These results coincides with the results obtained in EIT using non-stochastcic approaches based on modal expansion [33], the Weyl’s law [34], the classic near field-far field transformation [32], [14], and with the results based on field sampling theory [17], apart from a $\pi/4$ factor [^14] This discrepancy arises from the common sampling strategy in current electromagnetic field measurement systems, where sampling positions are constrained to lie along the coordinate curves of the scanning surface due to mechanical limitations [35]. [18].
@@
This method becomes an exact method to evaluate the $N_{DoF}\,\,$in the high-frequency asymptotic regime ($\lambda\rightarrow 0$), i.e., the geometrical optics approximation. In this regime, the $N_{DoF}\,\,$is strictly related to the étendue ($G_{e}$) [36], a pure geometrical quantity of great importance in radiometry [^15] Other interesting connections between radiometry and the stochastic approach used in this paper, that are not discussed here due to space limitations, are reported in [37]. [20](Sect. 5.7) since it is conserved under the geometrical optics approximation in the absence of diffraction, absorption, or scattering phenomena [^16] This invariance corresponds to the conservation of entropy in the absence of absorption, emission, or scattering processes [38]..
@@
Figure 3: From a ’spatial information’ point of view, linear source sees the scene in Fig. 2 in a distorted space.